OpenStudy (anonymous):

What is the solution of the equation? 4(y + 2) = 32 A. 4 B. 6 C. –10 D. 10

OpenStudy (ajprincess):

First remove the parentheses and then simplify.

OpenStudy (anonymous):

4y+8=32 4y=32-8 4y=24 y=24/4 y=6